# Hardness of Regular Expression Matching with Extensions

### v1

@@ 3 Preliminaries
$r::=a\mid\emptyset\mid r_{1}r_{2}\mid r_{1}+r_{2}\mid r^{*}$
where $a\in\Sigma$. For brevity, we use the character set notation $a_{1}a_{2}\cdots a_{k}]$, which is a shorthand for the regular expression $a_{1}+a_{2}+\cdots+a_{k}$ where $a_{1},\dots,a_{k}\in\Sigma$. In particular, we regard $\Sigma$ as the character set consisting of all elements of $\Sigma$. As standard, for a string or a regular expression $\alpha$ and a non-negative integer $i$, we write $\alpha^{i}$ for $\alpha$ concatenated $i$ times. The language of a regular expression is defined in the standard way. A regular expression $r$ matches a string $w$ if $w$ is an element of the language of $r$, which is also written as $w\in r$ by a slight abuse of notation.
4 OV to Semi-ERE Matching and Rewb Matching
In this section, we prove Theorems 1 and 2. The Orthogonal Vectors (OV) problem is the following: given sets $A,B\subseteq\{0,1\}^{d}$ containing the same number $n$ of bit-vectors of dimension $d$, decide if there are two vectors $a\in A$ and $b\in B$ such that $a$ and $b$ are orthogonal. A naive algorithm solves OV in $O(n^{2}d)$ time.
@@ 4 OV to Semi-ERE Matching and Rewb Matching
As mentioned in the introduction, it is known that SETH implies OVC (see Lemma A.1 in [47] for the proof), and therefore any hardness result under OVC also holds under SETH. Moreover, OVC is supported by evidence beyond its implication from SETH. See, e.g., [46] for more details.
@@ 4.1 Semi-ERE matching
For simplicity, we prove Theorem 2 before Theorem 1. Semi-ERE [23] extends regular expressions with the intersection construct $r_{1}\cap r_{2}$. The language of semi-ERE is defined by letting the language of $r_{1}\cap r_{2}$ be the intersection of the languages of $r_{1}$ and $r_{2}$. For example, the language of $(aa)^{*}\cap(aaa)^{*}$ is the same as that of $(aaaaaa)^{*}$, that is the set of strings $\{a^{6n}\mid n\geq 0\}$. The following lemma implies Theorem 2.
@@ 5 Clique to ERE Matching
In this section, we prove Theorem 3. ERE [23] extends regular expressions with the complement construct $\lnot r$. The language of $\lnot r$ is the complement of the language of $r$, that is, the set of strings over the alphabet that are not matched by $r$. For example, the language of $\lnot((aa)^{*})$ is the same as that of $a(aa)^{*}$ over the alphabet $\{a\}$. We use the intersection $r_{1}\cap r_{2}$ as a shorthand for $\lnot(\lnot r_{1}+\lnot r_{2})$.
The $k$-Clique problem asks whether a given undirected graph $G$ has a $k$-clique (i.e., the complete graph on $k$ vertices) as a subgraph. For a fixed $k$, a naive algorithm solves $k$-Clique in $O(n^{k})$ time. Abboud et al. have proposed a polylogarithmically faster combinatorial algorithm for the problem [4].
@@
Using fast matrix multiplication, Nešetřil and Poljak showed a faster algorithm that runs in $O(n^{\omega\lceil k/3\rceil})$ time [30]. Here, $2\leq\omega<2.3716$ [44] denotes the exponent of square matrix multiplication.
@@
We construct a reduction from $3k$-Clique to ERE matching as follows. Given a graph $G$ of $n$ vertices, we encode each vertex $v$ as a binary string $\bar{v}$ of fixed length $d=O(\log{n})$ and list all $k$-cliques $t_{1},\dots,t_{l}$ in $G$. We then encode them into a string $w$ as follows. For each $k$-clique $t=\{v_{1},\dots,v_{k}\}$, we define the following gadgets:
$VG(t)$ $=\bar{v_{1}}\textup{\%}\bar{v_{2}}\textup{\%}\cdots\bar{v_{k}}\textup{\%},$
$LG(t)$ $=\textup{\%}NG(v_{1})\textup{\%}NG(v_{2})\cdots\textup{\%}NG(v_{k}),$
$NG(v)$ $=\textup{\%}\bar{u_{1}}^{R}\textup{\%}\bar{u_{2}}^{R}\cdots\textup{\%}\bar{u_{p}}^{R}\quad\text{where $u_{1},\dots,u_{p}$ are the non-adjacent vertices of $v$}.$
For the sake of intuition, we name $VG(t)$ a vertex gadget, $LG(t)$ a list gadget, and $NG(v)$ a non-adjacent vertex gadget. Remark that a list gadget $LG(t)$ lists the non-adjacent vertices of each vertex of $t$. We then let
$w=\left(\bigodot_{i=1}^{l}\textup{\#}VG(t_{i})\right)\textup{\#}\textup{\textdollaroldstyle}\left(\bigodot_{i=1}^{l}\textup{\#}LG(t_{i})\textup{\textcentoldstyle}VG(t_{i})\right)\textup{\#}\textup{\textdollaroldstyle}\textup{\#}\left(\bigodot_{i=1}^{l}LG(t_{i})\textup{\#}\right).$
Both the length of $w$ and the time required for the construction are $O(k^{2}n^{k+1}\log{n})$.
@@
Let $t$ and $t^{\prime}$ be two $k$-cliques in $G$. Then, $r_{\mathrm{disconn}}$ matches $VG(t)\textup{\#}s\textup{\#}LG(t^{\prime})$ for some $s\in\Sigma^{*}$ if and only if $t\cup t^{\prime}$ is not a $2k$-clique in $G$.
@@
Suppose that $VG(t)\textup{\#}s\textup{\#}LG(t^{\prime})\in r_{\mathrm{disconn}}$. We can decompose it as $s_{1,0}s_{1,1}\textup{\#}s\textup{\#}s_{2,0}s_{2,1}$ where $VG(t)=s_{1,0}s_{1,1},LG(t^{\prime})=s_{2,0}s_{2,1}$ and $s_{1,1}\textup{\#}s\textup{\#}s_{2,0}\in r_{\mathrm{eq}}\cap[01]r_{\mathrm{eq}}[01]\cap\cdots\cap[01]^{d-1}r_{\mathrm{eq}}[01]^{d-1}$. From this and the definition of $r_{\mathrm{eq}}$, we observe that the middle substring $s_{1,1}\textup{\#}s\textup{\#}s_{2,0}$ has a binary encoding $\bar{v}$ of a vertex $v$ of $t$ as a prefix such that the reverse $\bar{v}^{R}$ is a member of $NG(v^{\prime})$ for some vertex $v^{\prime}$ of $t^{\prime}$. This implies that $v$ is not adjacent to $v^{\prime}$ and $t\cup t^{\prime}$ does not form a $2k$-clique in $G$. The proof of the other direction is analogous and omitted. ∎
@@
$w=w_{1,0}\textup{\#}VG(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}LG(t_{2})\textup{\textcentoldstyle}VG(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}LG(t_{3})\textup{\#}w_{3,1},$
where $w_{1,0},w_{3,1}\in[01\textup{\%}\textup{\#}]^{*}$ and $w_{1,1},w_{2,0},w_{2,1},w_{3,0}\in\Sigma^{*}$. Thus, it suffices to show that
  • $VG(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}LG(t_{2})\notin r_{\mathrm{disconn}}$,
  • $VG(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}LG(t_{3})\notin r_{\mathrm{disconn}}$ and
  • $VG(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}LG(t_{2})\textup{\textcentoldstyle}VG(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}LG(t_{3})\notin r_{\mathrm{disconn}}$,
@@
For the if direction, suppose that $r$ matches $w$. The delimiters # and \$ enforce that the subexpression $(\lnot r_{\mathrm{disconn}}\textup{\textcentoldstyle}\lnot r_{\mathrm{disconn}})\cap\lnot r_{\mathrm{disconn}}$ of $r$ matches a substring $s$ of $w$ of the form $s=VG(t_{1})\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\textdollaroldstyle}w_{3}\textup{\#}LG(t_{3})$ for some $k$-cliques $t_{1}$ and $t_{3}$ in $G$, and some strings $w_{1},w_{3}\in[01\textup{\%}\textup{\#}]^{*}$ and $w_{2}\in[01\textup{\%}\textup{\#}\textup{\textcentoldstyle}]^{*}$. From $s\notin r_{\mathrm{disconn}}$ and the previous claim, $t_{1}\cup t_{3}$ forms a $2k$-clique in $G$. Moreover, from $s\in\lnot r_{\mathrm{disconn}}\textup{\textcentoldstyle}\lnot r_{\mathrm{disconn}}$ and the use of the delimiter ¢, we can write $w_{2}$ as $w_{2,0}\textup{\#}LG(t_{2})\textup{\textcentoldstyle}VG(t_{2})\textup{\#}w_{2,1}$ for some $k$-clique $t_{2}$ in $G$ and strings $w_{2,0},w_{2,1}\in[01\textup{\%}\textup{\#}\textup{\textcentoldstyle}]^{*}$, and
  • $VG(t_{1})\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}LG(t_{2})\notin r_{\mathrm{disconn}}$ and
  • $VG(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3}\textup{\#}LG(t_{3})\notin r_{\mathrm{disconn}}$.
@@
$O(k^{2}n^{k+1}\log{n}+\log^{2}{n}+T(k^{2}n^{k+1}\log{n})\operatorname{polylog}(n))=\tilde{O}(T(n^{k+1})).\qed$
@@
We note that Lemma 9 also applies to Theorems 1 and 2. Therefore, unless OVC fails, neither 1-rewb matching nor semi-ERE matching can be solved within time both truly subquadratic in $n$ and constant-degree polynomial in $m$.

### v2

@@ 3 Preliminaries
$r::=a\mid\lambda\mid\emptyset\mid r_{1}r_{2}\mid r_{1}+r_{2}\mid r^{*}$
where $a\in\Sigma$ and $\lambda$ denotes the empty string. For brevity, we use the character set notation $a_{1}a_{2}\cdots a_{k}]$, which is a shorthand for the regular expression $a_{1}+a_{2}+\cdots+a_{k}$ where $a_{1},\dots,a_{k}\in\Sigma$. In particular, we regard $\Sigma$ as the character set consisting of all elements of $\Sigma$. As standard, for a string or a regular expression $\alpha$ and a non-negative integer $i$, we write $\alpha^{i}$ for $\alpha$ concatenated $i$ times. The language of a regular expression is defined in the standard way. A regular expression $r$ matches a string $w$ if $w$ is an element of the language of $r$, which is also written as $w\in r$ by a slight abuse of notation. Two regular expressions are equivalent if they define the same language.
4 OV to Semi-ERE, Rewb, and RSQ Matching
In this section, we prove Theorems 1, 2 and 3. The Orthogonal Vectors (OV) problem is the following: given sets $A,B\subseteq\{0,1\}^{d}$ containing the same number $n$ of bit-vectors of dimension $d$, decide if there are two vectors $a\in A$ and $b\in B$ such that $a$ and $b$ are orthogonal. A naive algorithm solves OV in $O(n^{2}d)$ time.
@@ 4 OV to Semi-ERE, Rewb, and RSQ Matching
As mentioned in the introduction, it is known that SETH implies OVC (see Lemma A.1 in [58] for the proof), and therefore any hardness result under OVC also holds under SETH. Moreover, OVC is supported by evidence beyond its implication from SETH. See, for example, [57] for more details.
@@ 4.1 Semi-ERE matching
For simplicity, we prove Theorem 3 before Theorems 1 and 2. Semi-ERE [29] extends regular expressions with the intersection construct $r_{1}\cap r_{2}$. The language of a semi-ERE is defined by letting the language of $r_{1}\cap r_{2}$ be the intersection of the languages of $r_{1}$ and $r_{2}$. For example, $(aa)^{*}\cap(aaa)^{*}$ is equivalent to $(aaaaaa)^{*}$, whose language is $\{a^{6n}\mid n\geq 0\}$. The following lemma implies Theorem 3.
@@ 4.2 Rewb matching
4.3 RSQ matching
In what follows, we prove Theorem 2 by establishing the following Lemma 8. RSQ [40] extends regular expressions with the squaring construct ${r}\{2\}$, which denotes an expression equivalent to $rr$.
Lemma 8.
If RSQ matching on a string of length $n$ and an RSQ of length $m$ can be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for some $\varepsilon>0$, then OV on sets of $n$ vectors of dimension $d$ can be solved in $O(n^{2-\varepsilon^{\prime}}\operatorname{poly}(d))$ time for some $\varepsilon^{\prime}>0$ and OVC is false.
Proof.
We construct a reduction from OV to RSQ matching as follows. Let $A$ and $B$ be sets of bit-vectors of dimension $d$ satisfying $|A|=|B|=n$. For $a\in A$, we write $a_{i}$ for the $i$-th coordinate of $a$ and let $\mathrm{VG}_{1}(a)=\textup{\textcentoldstyle}_{1}\overline{a_{1}}\textup{\textcentoldstyle}_{1}\overline{a_{2}}\cdots\textup{\textcentoldstyle}_{1}\overline{a_{d}}\textup{\textcentoldstyle}_{1}$ where $\overline{0}=1$ and $\overline{1}=11$ (e.g., $\mathrm{VG}_{1}(011)=\textup{\textcentoldstyle}_{1}1\textup{\textcentoldstyle}_{1}11\textup{\textcentoldstyle}_{1}11\textup{\textcentoldstyle}_{1}$). Similarly, let $\mathrm{VG}_{2}(b)=\textup{\textcentoldstyle}_{2}\overline{b_{1}}\textup{\textcentoldstyle}_{2}\overline{b_{2}}\cdots\textup{\textcentoldstyle}_{2}\overline{b_{d}}\textup{\textcentoldstyle}_{2}$ for $b\in B$. Let
$w=\left(\bigodot_{a\in A}\textup{\#}\mathrm{VG}_{1}(a)\right)\textup{\textdollaroldstyle}\left(\bigodot_{b\in B}\mathrm{VG}_{2}(b)^{R}\textup{\#}\right).$
Both the length of $w$ and the time required for this construction are $O(nd)$.
We then define an RSQ $r$ over the alphabet $\Sigma=\{1,\textup{\#},\textup{\textdollaroldstyle},\textup{\textcentoldstyle}_{1},\textup{\textcentoldstyle}_{2}\}$. We write $r?$ for $r+\lambda$, where $\lambda$ denotes the empty string. Let
$r$ $=[1\textup{\#}\textup{\textcentoldstyle}_{1}\textup{\textcentoldstyle}_{2}]^{*}\textup{\#}r_{d}\textup{\#}[1\textup{\#}\textup{\textcentoldstyle}_{1}\textup{\textcentoldstyle}_{2}]^{*},$
$r_{k}$ $=\textup{\textcentoldstyle}_{1}{(1(r_{k-1}?)(1?))}\{2\}\textup{\textcentoldstyle}_{2}\text{ for }k\in\{1,\dots,d\},$
$r_{0}$ $=\textup{\textcentoldstyle}_{1}\Sigma^{*}\textup{\textcentoldstyle}_{2}.$
Both the length of $r$ and the time required for this construction are $O(d)$. To establish the correctness of the reduction, it suffices to show the following claim:
Claim.
Let $a\in A$, $b\in B$ and $s$ be a string over $\Sigma$ where no substring of the form $\textup{\textcentoldstyle}_{2}\Sigma^{*}\textup{\textcentoldstyle}_{1}$ occurs. Then, $a$ and $b$ are orthogonal if and only if $r_{d}$ matches $\mathrm{VG}_{1}(a)s\mathrm{VG}_{2}(b)^{R}$.
Proof.
We prove by induction on $d$. The case $d=0$ is trivial. Suppose that $d\geq 1$ and $a$ and $b$ are orthogonal. Let $e=\textup{\textcentoldstyle}_{1}1r_{d-1}1\textup{\textcentoldstyle}_{2}\,+\,\textup{\textcentoldstyle}_{1}1r_{d-1}11\textup{\textcentoldstyle}_{2}\,+\,\textup{\textcentoldstyle}_{1}11r_{d-1}1\textup{\textcentoldstyle}_{2}$. Observe that $r_{d}$ is equivalent to
$\textup{\textcentoldstyle}_{1}(1(1?))^{2}\textup{\textcentoldstyle}_{2}\>+\>\textup{\textcentoldstyle}_{1}1r_{d-1}(1?)1r_{d-1}(1?)\textup{\textcentoldstyle}_{2}\>+\>\textup{\textcentoldstyle}_{1}(1r_{d-1}111\>+\>111r_{d-1}(1?)\>+\>11r_{d-1})\textup{\textcentoldstyle}_{2}+e$ ($\ast$)
and therefore $r_{d}$ matches strings that $e$ matches. We write $u$ for $\mathrm{VG}_{1}(a)s\mathrm{VG}_{2}(b)^{R}$ and $u^{\prime}$ for $\textup{\textcentoldstyle}_{1}\overline{a_{2}}\cdots\textup{\textcentoldstyle}_{1}\overline{a_{d}}\,\textup{\textcentoldstyle}_{1}\,s\,\textup{\textcentoldstyle}_{2}\,\overline{b_{d}}^{R}\textup{\textcentoldstyle}_{2}\cdots\overline{b_{2}}^{R}\textup{\textcentoldstyle}_{2}$. By the induction hypothesis, $r_{d-1}$ matches $u^{\prime}$. From $a_{1}b_{1}=0$, it follows that $u$ is one of $\textup{\textcentoldstyle}_{1}1u^{\prime}1\textup{\textcentoldstyle}_{2}$, $\textup{\textcentoldstyle}_{1}1u^{\prime}11\textup{\textcentoldstyle}_{2}$ and $\textup{\textcentoldstyle}_{1}11u^{\prime}1\textup{\textcentoldstyle}_{2}$. Hence $e$ matches $u$ and so does $r_{d}$.
Conversely, suppose that $r_{d}$ matches $u$. We claim that so does $e$. This implies that $(\overline{a_{1}},\overline{b_{1}})\in\{(1,1),(1,11),(11,1)\}$ and hence $a_{1}b_{1}=0$. This completes the proof, because $a_{i}b_{i}=0$ for $i\in\{2,\cdots,d\}$ by the induction hypothesis and thus $a$ and $b$ are orthogonal. It suffices to show that no term in ($\ast$4.3) other than $e$ matches $u$. It is straightforward to see that neither $\textup{\textcentoldstyle}_{1}(1(1?))^{2}\textup{\textcentoldstyle}_{2}$ nor $\textup{\textcentoldstyle}_{1}(1r_{d-1}111+111r_{d-1}(1?)+11r_{d-1})\textup{\textcentoldstyle}_{2}$ matches $u$. Moreover, if $\textup{\textcentoldstyle}_{1}1r_{d-1}(1?)1r_{d-1}(1?)\textup{\textcentoldstyle}_{2}$ were to match $u$, then $u$ would have a substring of the form $\textup{\textcentoldstyle}_{2}(1?)1\textup{\textcentoldstyle}_{1}$ because $r_{d-1}$ only matches strings that begin with $\textup{\textcentoldstyle}_{1}$ and end with $\textup{\textcentoldstyle}_{2}$. This contradicts the fact that $\mathrm{VG}_{1}(a)$ contains no $\textup{\textcentoldstyle}_{2}$, $\mathrm{VG}_{2}(b)$ contains no $\textup{\textcentoldstyle}_{1}$, and no substring of the form $\textup{\textcentoldstyle}_{2}\Sigma^{*}\textup{\textcentoldstyle}_{1}$ occurs in $s$. ∎
Without loss of generality, we may assume that $2-\varepsilon\geq 1$. Using the RSQ matching algorithm in the hypothesis, we can solve the OV problem with $d=c\log{n}$ for any constant $c\geq 1$ in time
$O(nd+d+(nd)^{2-\varepsilon}\operatorname{poly}(d))=O(n^{2-\varepsilon}\operatorname{poly}(d))=O(n^{2-\varepsilon/2}).\qed$
Remark 9.
As in Remark 6, we can avoid using the Kleene star in the construction of $r$ by adopting a technique used in Lemma 2 of [47]. Let $e^{*}$ be a starred subexpression of $r$. Instead of $e^{*}$, we use ${(\cdots({({e?}\{2\})}\{2\})\cdots)}\{2\}$ where squares are nested $\lceil\log_{2}{|w|}\rceil$ times. Then, $r$ matches $w$ if and only if it matches $w$ after this change, because the matching is determined by which subexpressions of $r$ match which substrings of $w$ and the lengths of the latter are at most $|w|$. Although both the length of $r$ and the time required for the construction increase to $O(d+\log{n})$, the resulting lower bound remains unchanged because we allow $d$ to be logarithmic in $n$. Thus, Lemma 8 holds even when restricted to star-free RSQs.
@@ 5 Clique to ERE Matching
In this section, we prove Theorem 4. ERE [29] extends regular expressions with the complement construct $\lnot r$. The language of $\lnot r$ is the complement of the language of $r$, that is, the set of strings over the alphabet that are not matched by $r$. For example, $\lnot((aa)^{*})$ is equivalent to $a(aa)^{*}$ over the alphabet $\{a\}$. We use the intersection $r_{1}\cap r_{2}$ as a shorthand for $\lnot(\lnot r_{1}+\lnot r_{2})$.
The $k$-Clique problem asks whether a given undirected graph $G$ has a $k$-clique (i.e., the complete graph on $k$ vertices) as a subgraph. For a fixed $k$, a naive algorithm solves $k$-Clique in $O(n^{k})$ time. Abboud et al. [4] have proposed a polylogarithmically faster combinatorial algorithm for the problem.
@@
Using fast matrix multiplication, Nešetřil and Poljak showed a faster algorithm that runs in $O(n^{\omega\lceil k/3\rceil})$ time [42]. Here, $2\leq\omega<2.3714$ [7] denotes the exponent of square matrix multiplication.
@@
We construct a reduction from $3k$-Clique to ERE matching as follows. Given a graph $G$ of $n$ vertices, we encode each vertex $v$ as a binary string $\overline{v}$ of fixed length $d=O(\log{n})$ and list all $k$-cliques $t_{1},\dots,t_{l}$ in $G$. We then encode them into a string $w$ as follows. For each $k$-clique $t=\{v_{1},\dots,v_{k}\}$, we define the following gadgets:
$\mathrm{VG}(t)$ $=\overline{v_{1}}\textup{\%}\overline{v_{2}}\textup{\%}\cdots\overline{v_{k}}\textup{\%},$
$\mathrm{LG}(t)$ $=\textup{\%}\mathrm{NG}(v_{1})\textup{\%}\mathrm{NG}(v_{2})\cdots\textup{\%}\mathrm{NG}(v_{k}),$
$\mathrm{NG}(v)$ $=\textup{\%}\overline{u_{1}}^{R}\textup{\%}\overline{u_{2}}^{R}\cdots\textup{\%}\overline{u_{p}}^{R}\quad\text{where $u_{1},\dots,u_{p}$ are the non-adjacent vertices of $v$}.$
For the sake of intuition, we name $\mathrm{VG}(t)$ a vertex gadget, $\mathrm{LG}(t)$ a list gadget, and $\mathrm{NG}(v)$ a non-adjacent vertex gadget. Remark that a list gadget $\mathrm{LG}(t)$ lists the non-adjacent vertices of each vertex of $t$. We then let
$w=\left(\bigodot_{i=1}^{l}\textup{\#}\mathrm{VG}(t_{i})\right)\textup{\#}\textup{\textdollaroldstyle}\left(\bigodot_{i=1}^{l}\textup{\#}\mathrm{LG}(t_{i})\textup{\textcentoldstyle}\mathrm{VG}(t_{i})\right)\textup{\#}\textup{\textdollaroldstyle}\left(\bigodot_{i=1}^{l}\textup{\#}\mathrm{LG}(t_{i})\right)\textup{\#}.$
Both the length of $w$ and the time required for the construction are $O(n^{k+1}\log{n})$.
@@
Let $t$ and $t^{\prime}$ be two $k$-cliques in $G$ and $s\in\Sigma^{*}$. Then, $r_{\mathrm{disconn}}$ matches $\mathrm{VG}(t)\textup{\#}s\textup{\#}\mathrm{LG}(t^{\prime})$ if and only if $t\cup t^{\prime}$ is not a $2k$-clique in $G$.
@@
Suppose that $\mathrm{VG}(t)\textup{\#}s\textup{\#}\mathrm{LG}(t^{\prime})\in r_{\mathrm{disconn}}$. We can decompose it as $s_{1,0}s_{1,1}\textup{\#}s\textup{\#}s_{2,0}s_{2,1}$ where $\mathrm{VG}(t)=s_{1,0}s_{1,1},\mathrm{LG}(t^{\prime})=s_{2,0}s_{2,1}$ and $s_{1,1}\textup{\#}s\textup{\#}s_{2,0}\in r_{\mathrm{eq}}\cap[01]r_{\mathrm{eq}}[01]\cap\cdots\cap[01]^{d-1}r_{\mathrm{eq}}[01]^{d-1}$. From this and the definition of $r_{\mathrm{eq}}$, we observe that the middle substring $s_{1,1}\textup{\#}s\textup{\#}s_{2,0}$ has a binary encoding $\overline{v}$ of a vertex $v$ of $t$ as a prefix such that the reverse $\overline{v}^{R}$ is a member of $\mathrm{NG}(v^{\prime})$ for some vertex $v^{\prime}$ of $t^{\prime}$. This implies that $v$ is not adjacent to $v^{\prime}$ and $t\cup t^{\prime}$ does not form a $2k$-clique in $G$. The proof of the other direction is analogous and omitted. ∎
@@
$w=w_{1,0}\textup{\#}\mathrm{VG}(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\textup{\textcentoldstyle}\mathrm{VG}(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}\mathrm{LG}(t_{3})\textup{\#}w_{3,1},$
where $w_{1,0},w_{1,1},w_{3,0},w_{3,1}\in[01\textup{\%}\textup{\#}]^{*}$ and $w_{2,0},w_{2,1}\in[01\textup{\%}\textup{\#}\textup{\textcentoldstyle}]^{*}$. Thus, it suffices to show that
  • $\mathrm{VG}(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\notin r_{\mathrm{disconn}}$,
  • $\mathrm{VG}(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}\mathrm{LG}(t_{3})\notin r_{\mathrm{disconn}}$ and
  • $\mathrm{VG}(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\textup{\textcentoldstyle}\mathrm{VG}(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}\mathrm{LG}(t_{3})\notin r_{\mathrm{disconn}}$,
@@
For the if direction, suppose that $r$ matches $w$. The delimiters # and \$ enforce that the subexpression $(\lnot r_{\mathrm{disconn}}\textup{\textcentoldstyle}\lnot r_{\mathrm{disconn}})\cap\lnot r_{\mathrm{disconn}}$ of $r$ matches a substring $s$ of $w$ of the form $s=\mathrm{VG}(t_{1})\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\textdollaroldstyle}w_{3}\textup{\#}\mathrm{LG}(t_{3})$ for some $k$-cliques $t_{1}$ and $t_{3}$ in $G$, and some strings $w_{1},w_{3}\in[01\textup{\%}\textup{\#}]^{*}$ and $w_{2}\in[01\textup{\%}\textup{\#}\textup{\textcentoldstyle}]^{*}$. From $s\notin r_{\mathrm{disconn}}$ and the previous claim, $t_{1}\cup t_{3}$ forms a $2k$-clique in $G$. Moreover, from $s\in\lnot r_{\mathrm{disconn}}\textup{\textcentoldstyle}\lnot r_{\mathrm{disconn}}$ and the use of the delimiter ¢, we can write $w_{2}$ as $w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\textup{\textcentoldstyle}\mathrm{VG}(t_{2})\textup{\#}w_{2,1}$ for some $k$-clique $t_{2}$ in $G$ and strings $w_{2,0},w_{2,1}\in[01\textup{\%}\textup{\#}\textup{\textcentoldstyle}]^{*}$, and
  • $\mathrm{VG}(t_{1})\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\notin r_{\mathrm{disconn}}$ and
  • $\mathrm{VG}(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3}\textup{\#}\mathrm{LG}(t_{3})\notin r_{\mathrm{disconn}}$.
@@
$O(n^{k+1}\log{n}+\log^{2}{n}+T(n^{k+1}\log{n})\operatorname{poly}(\log^{2}{n}))=\tilde{O}(T(n^{k+1})).\qed$
Remark 11.
[…]
@@
We note that Lemma 14 also applies to Theorems 1, 2 and 3. Therefore, unless OVC fails, neither 1-rewb matching, RSQ matching nor semi-ERE matching can be solved within time both truly subquadratic in $n$ and constant-degree polynomial in $m$.